

********************************************************************
UARS WEEKLY STATUS REPORT - 26 Apr 2002
********************************************************************
***********
SOLSTICE
***********
Nothing received.
***********
SUSIM
***********
SUSIM began its first scans after yaw on April 17. Since then,
the instrument has performed its daily mid-resolution and MgII
solar scans during periods of predicted UARS telemetry.
Ed Esfandiari has written IDL routines to automatically generate
(on the fly) the RSS mode DCL runsteams needed for SUSIM
reprocessing. Using these tools, we were able to reprocess at
level-1 more than 1000 days of SUSIM UARS data in just 2 days! We
then estimate that we will be able reprocess the entire SUSIM
dataset from a low level 1 to level 3BS in about a week.
***********
ACRIM
***********
Nothing received.
***********
PEM
***********
PEM continues to receive NMC data files from NOAA. No problems have
been encountered.
PEM support of the UARS Traceability Mission for the TIMED mission has
been organized with the GVUI team. Predicted TIMED-UARS conjunctions from
the high geographic latitudes were compared with the expected magnetic position
and the PEM climatology. Positional and anticipated precipitation flux
information showed that the expected electron precipitation as a function of
magnetic activity had the highest probably of a successful conjunction on
April 19. Continued monitoring has shown changes in conjunction times. For
conjunctions within 300 km, there are now 4 times shown:
Contact # Date Day Time Geo Lat ILAT MLT Hemisphere
(deg) (deg) (hr)
1 April 18 108 21:30 56.87 54.43 2.96 North
2 April 18 108 22:18 -56.89 66.06 6.07 South
3 April 23 113 14:56 56.98 53.06 20.04 North
4 April 23 113 15:44 -56.98 47.76 9.11 South
Contact times have remained about the same in geographic position, but not in
magnetic position. The MLT of contact 3 has changed the most. Definitive UARS
orbit positions will determine the final locations.
*Beginning of power orbit day (PEM activation): array current >5 amps, array
voltage >53.5 volts, and battery current >2.0 amps.
End of power orbit day (PEM deactivation): time in power orbit day >4 minutes,
array current <9 amps, and battery currents <0.2 amps.
Power orbit day variation: ~39-40 min at yaw to ~50-51 min at maximum beta.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
Nothing received.
***********
WINDII
***********
Nothing received.
***********
MPG
***********
The MPG continues to serve as the point-of-contact
for UARS instrument operations to support re-engineering
of the UARS Mission Operations Concept to develop
capabilities for the Science Traceability Mission.
Daily processing of instrument Daily Activity Plans
continues with no problems.
***********
CDHF
***********
Nothing received.
***********
FOT
***********
REPORTING PERIOD
This report covers 13 April 2002 (Orbit 57889, GMT Day 103) through 19 April
2002 (Orbit 57994, GMT Day 109).
SPACECRAFT OPERATION
The observatory is now in Forward Flight (Southern Hemisphere viewing)
and is performing nominally under two-battery operations. The
instrument operational changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
PEM 106/1340 57943 Power ON
SUSIM 107/1338 57958 Power ON START
SUSIM 107/1437 57958 Power ON CONTINUED
SUSIM 108/1254 57977 Power ON COMPLETED
PEM 108/1336 57871 Power ON
HALOE 109/1436 57988 Power ON
HALOE 109/2011 57992 Power ON COMPLETED
A Reverse-to-Forward Yaw Around maneuver (number 108) was successfully
executed in Orbit 57899 beginning 103/16:10:22Z and completing at
103/16:47:32Z on 13 April 2002. The SOLSTICE instrument was in day
only and all other instruments were in their off or standby mode
(including HRDI and PEM). Due to the SA position at 269 degrees, the
effective night length of the spacecraft after the maneuver was
approximately 56.1 minutes. Post maneuver analysis of data indicate
that during the Yaw Around, the end-of-night load bus voltage was
24.30 volts and the average depths of discharge (DODs) on Batteries 2
and 3 were approximately 21.09 and 20.62 percent respectively. The
next (forward-to-reverse) YAW is scheduled for 23 May 2002 (DOY 143).
The beta angle ranged between the angles of 1.58 degrees to 25.85
degrees for this report period. The beta angle is now increasing
towards a maximum angle of 73 degrees on 03 May 2002 (DOY 123).
Spacecraft battery 1, 2, and 3 performance monitoring continues.
Battery 1 remains OFF the charge relay. Battery 2 and Battery 3 are
currently maintaining greater than 23.8 V end-of-night (EON) load bus
voltage. Battery temperatures are stable with a temperature delta
between Battery 2 and Battery 3 of 0.96 to 1.59 degrees C. Additional
battery performance data for each battery on 12 April 2002 and 19
April 2002 are provided in Appendix D.
The clock error ranged from -14.0 to 13.1 msec during this report
period. Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial
excursion of ALERTS or Out-Of-Limits conditions detected by the FOT
for this report. A summary is listed below and the details of ALERT
and Out-Of-Limits occurrences are listed in Appendix B.
OUT-OF-LIMITS = 00
ALERTS = 00
The following real-time command sheets were executed for the listed
operational element(s) during this report period (see Appendix A).
FOT = 09
HALOE = 02
PEM AXIS = 02
SUSIM = 03
UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.
FOT = 1
Three (3) Anomaly Reports remained OPEN at the end of this report
period (see Appendix E).
TRANSPONDER FREQUENCIES:
Transponder A: 01/333 - 2287.496855
Transponder B: 02/043 - 2287.498879
GROUND SYSTEM OPERATION
One (1) operational support problems occurred this period. Anomaly
Reports are included in Appendix E.
The data loss calculations are no longer being supplied by Data
Capture Facility (DCF).
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
The FOT successfully executed the yaw maneuver on 13 April 2002 from
the new Eclipse ground system. There were no problems encountered
with the new system during the maneuver. The instrument power offs
and turn ons went well also with only a few minor timing problems
which can easily be remedied before the next instrument configuration
change. The Eclipse system appears to have passed its test during the
maneuver period and only a few minor modifications are left to be
completed. Once again the FOT would like to thank all those who had a
part in implementing the new ground system.
MEETINGS
The FOT did not attended any meetings this week.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw slew is scheduled for 23 May 2002 (DOY 143). This
maneuver will be a Forward to Reverse Flight maneuver (Northern
Hemisphere viewing).
Attachments: Appendix A, Real-time Command Sheets
Appendix B, Out-Of-Limits & ALERT
Occurrences
Appendix C, Clock & Solar Array Rate
Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 19 April 2002
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ ------
57896 103/1136 02-126 OBC/CLK CLK ADJ FR 20.6 TO 0.0 MSEC/DAY
57899 103/1539 02-127 MACS TAM TO MONITER MODE
57899 103/1539 02-128 MACS/OBC YAW SCRIPT
57899 103/1627 02-129 MACS TAM ALIGN/UPDATE MODE
57901 103/1851 02-130 OBC/CLK CLK ADJ FR 0.0 TO -20.6 MSEC/DAY
57925 105/0939 02-131 OBC/CLK CLK ADJ FR -20.6 TO 0.0 MSEC/DAY
57932 105/1958 02-132 OBC/CLK CLK ADJ FR 0.0 TO 20.6 MSEC/DAY
57943 106/1340 02-133 PEM AXIS POWER ON
57947 106/1846 02-134 OBC/CLK CLK ADJ FR 20.6 TO 0.0 MSEC/DAY
57956 107/1018 02-135 OBC/CLK CLK ADJ FR 0.0 TO -20.6 MSEC/DAY
57958 107/1338 02-136 SUSIM POWER ON START
57958 107/1437 02-136 SUSIM POWER ON CONTINUED
57977 108/1254 02-136 SUSIM POWER ON COMPLETED
57978 108/1336 02-137 PEM POWER ON
57988 109/1436 02-139 HALOE POWER ON
57992 109/2011 02-140 HALOE POWER ON COMPLETED
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 19 April 2002
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
-------- ----- ----------- ------
-None this report period.
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
-------- ----------- ------
-None this report period.
______________________________________________________________________
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 19 April 2002
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ----- -----------------
57896 103/1140 CLK ADJ FR 20.6 TO 0.0 MSEC/DAY
57901 103/1839 CLK ADJ FR 0.0 TO -20.6 MSEC/DAY
57925 105/0945 CLK ADJ FR -20.6 TO 0.0 MSEC/DAY
57932 105/2010 CLK ADJ FR 0.0 TO 20.6 MSEC/DAY
57947 106/1846 CLK ADJ FR 20.6 TO 0.0 MSEC/DAY
57956 107/1023 CLK ADJ FR 0.0 TO -20.6 MSEC/DAY
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 19 April 2002
12 April 2002 - GMT Day 102 Beta = 7.14 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 +728.0/+728.0 -3.16/-4.38 +0.0/ +0.0 0.0 0.000 0.0
2 +11.2/-134.0 +4.50/+3.87 +27.8/-15.8 25.0 1.010 20.6
3 +22.4/-22.4 +3.55/+2.60 +26.7/-15.8 25.0 1.010 20.3
19 April 2002 - GMT Day 109 Beta = 25.85 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 +728.0/+728.0 -3.16/-4.38 +0.0/ +0.0 0.0 0.000 0.0
2 +11.2/ -89.6 +5.14/+4.18 +29.2/-15.2 24.8 1.030 20.7
3 +22.4/-33.6 +4.50/+2.91 +28.4/-15.2 24.8 1.020 20.4
V/T LEVEL CHANGES
ORB# TIME LVL FR-TO
----- ---- ---------
-no changes V/T 5
APPENDIX E
UARS ANOMALY REPORTS
Week ending 19 April 2002
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
02-025 FOT 57899 103/1627 14170 Generic Late Acquisition
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
02-025 FOT See List Apr 13 N/A Generic Late Acquisition
ANOMALY INVESTIGATION REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period.
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
02-004 SOLSTICE 56883 036/0050 N/A Solstice in Wrong Mode
02-023 FOT 57767 094/2027 24285 Early LOS on OMNI Pass
02-024 FOT 57783 095/2209 24285 SOLSTICE INOPS in Day